OpenStudy (anonymous):

What is limit of x * ln(x) when x approaches 0. Please prove it.

OpenStudy (anonymous):

Someone answered x * ln(x) = ln(x) / (1 / x). Using L'Hospital I get that the limit is 0. Am I right?

OpenStudy (anonymous):

Actually it should be \[\large \frac{x}{\frac{1}{\ln x}} \]Because this is an actual indeterminate form.

OpenStudy (anonymous):

Look, since \(\ln(x)\) isn't defined for negative values, then the limit can't exist, right?

OpenStudy (anonymous):

You can really only do a left hand limit. Is that your goal?